# Spending Behavior and Economic Impacts of Urban Digital Consumption Vouchers

### v1

@@ 1 Introduction
Several countries implemented large-scale fiscal interventions in response to the COVID-19 pandemic. In the United States, the Coronavirus Aid, Relief, and Economic Security (CARES) Act, enacted in 2020, amounted to a USD 2.2 trillion stimulus package, which was the largest in U.S. history. It included direct cash payments to individuals, expanded unemployment insurance, and sector-specific financial support. Japan introduced three rounds of stimulus packages totaling more than 300 trillion yen, equivalent to approximately 60 percent of its GDP. These measures featured universal cash payments along with subsidies for fuel and utility expenses. South Korea also provided direct cash payments to all households, and Singapore launched the Care and Support Package to deliver one-time cash transfers. In contrast, Taiwan not only launched general-purpose consumption vouchers, as in other countries, but also adopted a more targeted strategy by issuing vouchers with restricted usage conditions, aiming to encourage spending in specific sectors that were most severely affected by the pandemic.
@@
In contrast, consumption vouchers may help preserve the intended stimulative effect by limiting flexibility and channeling spending toward targeted sectors. While Kan et al. (2017) reported limited effects from Taiwan’s 2008 paper-based voucher program, Xing et al. (2023) and Chen et al. (2025) documented strong stimulus outcomes associated with digital consumption vouchers in China. Moreover, the latter two studies highlight several advantages of digital voucher schemes, including lower administrative costs, faster distribution, real-time tracking of transactions, and the ability to support specific industries such as retail and food services.
Nevertheless, the existing literature on voucher usage and effectiveness faces two main limitations. First, most studies rely on telephone surveys, which make it difficult to verify the authenticity of responses and whether the vouchers were actually used (e.g., Kan et al. (2017)). Second, much of the literature focuses on the effects of a single voucher type, without investigating how different consumer groups respond to various types of vouchers (e.g., Leone and Srinivasan (1996); Xing et al. (2023); Chen et al. (2025)).
@@
This paper reports three main findings. First, expenditure substitution rates vary significantly across voucher types. Sports vouchers exhibit the highest substitution rates, suggesting that spending in this category often replaces pre-planned consumption, likely due to the habitual nature of sports-related purchases. In contrast, accommodation vouchers show the lowest substitution rates, indicating greater potential to generate new spending. Second, induced consumption effects are substantial. The accommodation voucher again stands out, with the highest induced consumption, reinforcing its effectiveness in promoting incremental and geographically distributed spending. Third, regional input–output analysis reveals that the baseline output multiplier of the Bear Vouchers 2.0 program is 0.97 when behavioral responses are not considered. Once substitution and induced consumption effects are incorporated, the multiplier rises to as high as 1.76, indicating that each NT\$1 in voucher spending can generate up to NT\$1.76 in regional economic output. These findings highlight the critical role of consumer behavior in shaping the effectiveness of voucher-based fiscal stimulus and underscore the importance of careful program design tailored to spending elasticity and sectoral targeting. Analysis of treatment intensity further shows that even modest increases in voucher face value can generate meaningful marginal spending effects, especially when such increases are unexpected or paired with vendor promotions.
@@ 2.2 Consumption Vouchers
Building on these insights,Chen et al. (2025) evaluated a program in Ningbo, China, and found that digital coupons significantly raised restaurant revenues, with a return of 4.5 yuan per 1 yuan of government spending. However, the effects were temporary, and while large businesses captured most absolute gains, smaller businesses experienced greater relative growth, contributing to reduced revenue and welfare inequality. These findings suggest that digital coupons are effective short-term support tools but are better suited for temporary relief rather than structural economic reform.
@@ 3 Background and Data
In addition, the Taipei City Government introduced an extra round of the voucher program on December 16 of the same year. All individuals who had previously registered for the Bear Vouchers 2.0 lottery were eligible to receive additional vouchers corresponding to the types they originally selected, regardless of whether they had won in the first round. The values of these extra vouchers varied by type: NT\$500 for accommodation, and NT\$100 each for dining, cultural, sports, market, and agricultural vouchers. For instance, a participant who had registered for both accommodation and dining vouchers would receive one additional voucher of each type, totaling NT\$600 in value.
The dataset employed in this paper is derived from a survey conducted among individuals who actually utilized the Taipei Bear Vouchers 2.0. The survey was administered between March 1 and March 8, 2023, yielding a total of 159,211 valid responses. Based on the survey results, this paper estimates consumer behavioral parameters for each type of digital voucher and are detailed in the following section.
@@ 4.2 Induced Consumption
Responses to this question were categorized into different spending bands depending on the type of voucher. The details of these brackets can be found in the Appendix. Using these intervals, the induced consumption rate for subgroup $j$ receiving voucher type $k$ is computed as:
$IC_{jk}=\sum_{c=1}^{C}\frac{m_{kc}\times b_{jkc}}{F_{k}},$ (3)
where $C$ denotes the total number of expenditure intervals and $F_{k}$ represents the face value of voucher type $k$. The term $m_{kc}$ indicates the midpoint of the $c$-th interval, with the convention that $m_{kc}=0$ when the interval corresponds to no additional spending, and that for the highest interval, $m_{kc}$ is set to the lower bound of that uppermost range. Additionally, we define $b_{jkc}=\frac{1}{n_{j,k}}\sum_{i=1}^{n_{j,k}}\mathbf{1}_{AS_{i}\in c}$, where $AS_{i}$ refers to the out-of-pocket additional spending reported by individual $i$ in subgroup $j$, and $\mathbf{1}_{AS_{i}\in c}$ is an indicator function that equals 1 if $AS_{i}$ falls within interval $c$, and 0 otherwise. As such, the product $m_{kc}\times b_{jkc}$ reflects the average additional spending associated with interval $c$, and dividing this by $F_{k}$ expresses the amount relative to the voucher’s face value. A value of $IC_{jk}>1$ indicates that, on average, respondents’ extra spending surpassed the voucher amount, suggesting a strong induced consumption effect on total consumption. Similarly, we can define overall induced consumption rate for whole respondents who received voucher type $k$:
$IC_{k}=\sum_{c=1}^{C}\frac{m_{kc}\times b_{kc}}{F_{k}},$ (4)
where $b_{kc}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\mathbf{1}_{AS_{i}\in c}$.
@@ 4.3 Intensity of Treatment
$IT_{k}=\sum_{c=1}^{C}m_{kc}\times(b_{kc,1}-b_{kc,2}),$ (5)
where $b_{kc,1}$ and $b_{kc,2}$ are the proportions of respondents using original and additional stimulus vouchers whose reported additional spending fall into interval $c$, respectively. [^1] The calculation of $b_{kc,1}$ and $b_{kc,2}$ are similar to those defined previously. The difference $b_{kc,1}-b_{kc,2}$ reflects the marginal effect of moving from lower- to higher-intensity treatment for each spending bracket, and it delivers different meanings compared with the induced income rate.
@@ 4.4 Identification and Inference
where ${\widehat{y}}_{jk}$ could be either $ES_{jk}$ or $IC_{jk}$, and ${\widehat{B}}_{k}=\min_{j}{\widehat{y}}_{jk}$ represents the most conservative estimate of the bias. On this basis, we derive an interval estimate for the true effect $\theta_{jk}$, where the lower bound is adjusted to account for potential bias, and the upper bound corresponds to the unadjusted sample mean. This correction assumes that the smallest observed group mean provides a uniform upper bound on the reporting bias:
@@
To quantify the uncertainty of these estimates, we implement a stratified bootstrap procedure to construct $100(1-\alpha)\%$ confidence intervals for both bounds as shown in Algorithm 1.
@@ 5.2 Expenditure Substitution Effect
5.3 Induced Consumption Effect
5.4 Intensity of Treatment Effect
@@
(a) Accommodation
(b) Dining
(c) Cultural
(d) Sports
(e) Market
(f) Agricultural
Figure 2: Expenditure Substitution Rate
@@ 5.5 Economic Impacts of Consumption Vouchers
To assess the economic impact of Taipei’s consumption voucher program, we apply the Taipei Input–Output model to evaluate the effects of the Bear Vouchers 2.0 policy and to calculate the associated output multipliers based on Equation (8). In this framework, the estimated total induced consumption is treated as a change in final demand, denoted by $\Delta\mathbf{F}$. Specifically, for each voucher type $k$, we compute the induced consumption as the product of the original issued amount and the behavioral adjustment factor $(1-ES_{k})\times(1+IC_{k})$, where $ES_{k}$ denotes the expenditure substitution rate and $IC_{k}$ represents the induced consumption rate. These adjusted values are then used as inputs in the input–output model to estimate the broader economic impacts. The coefficient matrix and value-added vector used in Equation (8) are provided in Table 7 in the Appendix.
@@
Table 5: Estimated Policy Inputs for the Taipei’s Bear Vouchers 2.0 Program
@@
Note: The original issued amounts are provided by the Taipei City Government, and these figures include the second-round extra stimulus vouchers. The optimistic and pessimistic estimates of the induced change in demand are calculated using the expression $(1-ES_{k})\times(1+IC_{k})$, where $ES_{k}$ denotes the substitution effect and $IC_{k}$ denotes the induced consumption effect for voucher type $k$. These estimates are obtained by multiplying the issued amount by the above expression, using the mean values from the lower and upper bounds of the respective confidence intervals for substitution and induced consumption effects.
Table 6 presents the estimated economic benefits of the Bear Vouchers 2.0 program on Taipei City’s industrial economy under three scenarios. The baseline scenario reflects the economic impact based solely on the original budget allocation, treating it as a direct change in final demand without accounting for expenditure substitution or induced consumption effects. This scenario assumes a total policy input of NT\$584.53 million. In addition to the baseline, the table includes a pessimistic scenario and an optimistic scenario, both of which incorporate behaviorally adjusted estimates of final demand based on substitution and induced consumption effects as shown in Table 5.
@@
Table 6: Economic Benefit of Taipei’s Bear Vouchers 2.0 Program
@@
Note: The output multiplier is defined as the ratio of the change in gross domestic product (GDP) to the original amount of policy expenditure. The original amount of policy expenditure is NT\$584.53 million. All units are in measured in million except the output multiplier.

### v2

@@ 1 Introduction
Several countries implemented large-scale fiscal interventions in response to the COVID-19 pandemic. In the United States, the Coronavirus Aid, Relief, and Economic Security (CARES) Act, enacted in 2020, amounted to a USD 2.2 trillion stimulus package, which was the largest in U.S. history. It included direct cash payments to individuals, expanded unemployment insurance, and sector-specific financial support. Japan introduced three rounds of stimulus packages totaling more than 300 trillion yen, equivalent to approximately 60 percent of its GDP. These measures featured universal cash payments along with subsidies for fuel and utility expenses. South Korea also provided direct cash payments to all households, and Singapore launched the Care and Support Package to deliver one-time cash transfers. In contrast, Taiwan not only launched general-purpose consumption vouchers, as in other countries, but also adopted a more targeted strategy by issuing vouchers with restricted usage conditions, aiming to encourage spending in specific sectors that were most severely affected by the pandemic. [^1] In Taiwan, the government also introduced a relief and revitalization plan for small and medium-sized enterprises, the effects of which have been analyzed in Chen et al. (2025a).
@@
In contrast, consumption vouchers may help preserve the intended stimulative effect by limiting flexibility and channeling spending toward targeted sectors. While Kan et al. (2017) reported limited effects from Taiwan’s 2008 paper-based voucher program, Xing et al. (2023) and Chen et al. (2025b) documented strong stimulus outcomes associated with digital consumption vouchers in China. Moreover, the latter two studies highlight several advantages of digital voucher schemes, including lower administrative costs, faster distribution, real-time tracking of transactions, and the ability to support specific industries such as retail and food services.
Nevertheless, the existing literature on voucher usage and effectiveness faces two main limitations. First, most studies rely on telephone surveys, which make it difficult to verify the authenticity of responses and whether the vouchers were actually used (e.g., Kan et al. (2017)). Second, much of the literature focuses on the effects of a single voucher type, without investigating how different consumer groups respond to various types of vouchers (e.g., Leone and Srinivasan (1996); Xing et al. (2023); Chen et al. (2025b)).
@@
This paper reports three main findings. First, expenditure substitution rates vary significantly across voucher types. Sports vouchers exhibit the highest substitution rates, suggesting that spending in this category often replaces pre-planned consumption, likely due to the habitual nature of sports-related purchases. In contrast, accommodation vouchers show the lowest substitution rates, indicating greater potential to generate new spending. Second, induced consumption effects are substantial. The accommodation voucher again stands out, with the highest induced consumption, reinforcing its effectiveness in promoting incremental and geographically distributed spending. Third, regional input–output analysis reveals that the baseline output multiplier of the Taipei Bear Vouchers 2.0 program is 0.97 when behavioral responses are not considered. Once substitution and induced consumption effects are incorporated, the multiplier rises to as high as 1.76, indicating that each NT\$1 in voucher spending can generate up to NT\$1.76 in regional economic output. These findings highlight the critical role of consumer behavior in shaping the effectiveness of voucher-based fiscal stimulus and underscore the importance of careful program design tailored to spending elasticity and sectoral targeting. Analysis of treatment intensity further shows that even modest increases in voucher face value can generate meaningful marginal spending effects, especially when such increases are unexpected or paired with vendor promotions.
@@ 2.2 Consumption Vouchers
Building on these insights, Chen et al. (2025b) evaluated a program in Ningbo, China, and found that digital coupons significantly raised restaurant revenues, with a return of 4.5 yuan per 1 yuan of government spending. However, the effects were temporary, and while large businesses captured most absolute gains, smaller businesses experienced greater relative growth, contributing to reduced revenue and welfare inequality. These findings suggest that digital coupons are effective short-term support tools but are better suited for temporary relief rather than structural economic reform.
@@ 3 Background and Data
In addition, the Taipei City Government introduced an extra round of the voucher program on December 16 of the same year. All individuals who had previously registered for the Taipei Bear Vouchers 2.0 lottery were eligible to receive additional vouchers corresponding to the types they originally selected, regardless of whether they had won in the first round. The values of these extra vouchers varied by type: NT\$500 for accommodation, and NT\$100 each for dining, cultural, sports, market, and agricultural vouchers. For instance, a participant who had registered for both accommodation and dining vouchers would receive one additional voucher of each type, totaling NT\$600 in value.
The dataset employed in this paper is derived from a survey conducted among individuals who actually utilized the Taipei Bear Vouchers 2.0. The survey was administered between March 1 and March 8, 2023, yielding a total of 159,211 valid responses. Based on the survey results, this paper estimates consumer behavioral parameters for each type of digital voucher and are detailed in the following section. [^2] To enhance the response rate, we offered a NT\$50 digital coupon (Family Mart) as a lottery incentive, which, according to prior studies (e.g., Bosnjak and Tuten (2003); Sauermann and Roach (2013)), does not compromise response quality when used at a moderate level.
@@ 4.2 Induced Consumption
Responses to this question were categorized into discrete spending intervals, depending on the type of voucher received. The details of these brackets are provided in the Appendix. Using these intervals, the induced consumption rate for subgroup $j$ receiving voucher type $k$ is computed as:
$IC_{jk}=\sum_{c=1}^{C}\frac{m_{kc}\times b_{jkc,\mathrm{orig}}}{F_{k}^{\mathrm{orig}}}+\sum_{c=1}^{C}\frac{m_{kc}\times b_{jkc,\mathrm{bonus}}}{F_{k}^{\mathrm{bonus}}},$ (3)
where $C$ denotes the total number of expenditure intervals, and $F_{k}^{l}$ represents the face value of voucher type $k$ from either the original or bonus round, with $l\in\{\mathrm{orig},\mathrm{bonus}\}$. The term $m_{kc}$ denotes the midpoint of interval $c$; we follow the convention that $m_{kc}=0$ for the interval indicating no additional spending, and for the highest interval, $m_{kc}$ is set to the lower bound of that uppermost range. We define:
$b_{jkc,l}=\frac{1}{n_{j,k}}\sum_{i=1}^{n_{j,k}}\mathbf{1}_{AS_{i}\in c\cap F_{ki}=F_{k}^{l}},$
where $AS_{i}$ refers to the out-of-pocket additional spending reported by individual $i$ in subgroup $j$; $F_{ki}$ denotes the face value of voucher type $k$ received by individual $i$, which can be either $F_{k}^{\mathrm{orig}}$ or $F_{k}^{\mathrm{bonus}}$, depending on the round of issuance; and the indicator function $\mathbf{1}_{AS_{i}\in c\cap F_{ki}=F_{k}^{l}}$ equals 1 if $AS_{i}$ falls in interval $c$ and the face value of the voucher received by individual $i$ is $F_{k}^{l}$, and 0 otherwise.
Thus, the product $m_{kc}\times b_{jkc,l}$ represents the average additional spending within interval $c$, and dividing by $F_{k}^{l}$ scales this amount relative to the corresponding voucher face value. A value of $IC_{jk}>1$ implies that, on average, respondents in subgroup $j$ spent more out-of-pocket than the voucher amount, suggesting a strong induced consumption effect. Similarly, we define the overall induced consumption rate among all respondents who received voucher type $k$ as:
$IC_{k}=\sum_{c=1}^{C}\frac{m_{kc}\times b_{kc,\mathrm{orig}}}{F_{k}^{\mathrm{orig}}}+\sum_{c=1}^{C}\frac{m_{kc}\times b_{kc,\mathrm{bonus}}}{F_{k}^{\mathrm{bonus}}},$ (4)
where
$b_{kc,l}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\mathbf{1}_{AS_{i}\in c\cap F_{ki}=F_{k}^{l}}.$
@@ 4.3 Intensity of Treatment
$IT_{k}=\sum_{c=1}^{C}m_{kc}\times(b_{kc,\mathrm{orig}}-b_{kc,\mathrm{bonus}}),$ (5)
where $b_{kc,\mathrm{orig}}$ and $b_{kc,\mathrm{bonus}}$ are the proportions of respondents using original and additional stimulus vouchers whose reported additional spending fall into interval $c$, respectively. The difference $b_{kc,\mathrm{orig}}-b_{kc,\mathrm{bonus}}$ reflects the marginal effect of moving from lower- to higher-intensity treatment for each spending bracket, and it delivers different meanings compared with the induced income rate.
@@ 4.4 Identification and Inference
where ${\widehat{y}}_{jk}$ could be either $ES_{jk}$ or $IC_{jk}$, and ${\widehat{B}}_{k}=\min_{j}{\widehat{y}}_{jk}$ represents the most conservative estimate of the bias. On this basis, we derive an interval estimate for the true effect $\theta_{jk}$, where the lower bound is adjusted to account for potential bias, and the upper bound corresponds to the unadjusted sample mean. This correction strategy does not assert that the true effect in the lowest observed subgroup is literally zero. Rather, it adopts a worst-case bounding perspective, treating the minimum observed outcome as an upper bound on potential bias. That is,
@@
To quantify the uncertainty of these estimates, we implement a stratified bootstrap procedure to construct $100(1-\alpha)\%$ confidence intervals for both bounds as shown in Algorithm 1.333Formally, the minimum operator is not fully differentiable only when multiple subgroup estimates tie exactly. Given that self-reporting bias is expected to vary across groups, such ties occur with probability zero asymptotically. Hence, the condition in Fang and Santos (2019, Thm 3.1) for bootstrap consistency is satisfied.
@@ 4.5 Regional Input–Output Analysis
In addition, since Taiwan’s official statistics provide input–output tables only at the national level, this paper constructs a Taipei City input–output table by combining the national input–output table with regional statistics from Taiwan’s Industry and Service Census Report, including data on output, value added. This regional table serves as the basis for subsequent analysis. For details on the methodology and technical procedures used to compile the Taipei City input–output table, please refer to the Appendix.
@@ 5.2 Expenditure Substitution Effect
Compared to the substitution rates observed under the Taipei Bear Vouchers 1.0 program (Taipei City Government, 2022), the accommodation voucher also recorded the lowest rate (approximately 8%), while the sports voucher exhibited the highest (around 39%). The substitution rates for the dining, market, and cultural vouchers ranged between 24% and 30%, indicating a degree of behavioral inertia among consumers in their use of consumption vouchers. Notably, the substitution rates for the accommodation and sports vouchers were lower in the Taipei Bear Vouchers 1.0 program than in Taipei Bear Vouchers 2.0, whereas the other voucher types displayed comparatively lower substitution rates in the Taipei Bear Vouchers 2.0 program. [^4] Taipei City Government (2022) does not report substitution rate data for the Taipei Bear Vouchers 1.0 program disaggregated by demographic characteristics.
@@
(a) Accommodation
(b) Dining
(c) Cultural
(d) Sports
(e) Market
(f) Agricultural
Figure 1: Expenditure Substitution Rate
@@
5.3 Induced Consumption Effect
As an additional note, the induced consumption effect observed under the Taipei Bear Vouchers 2.0 program differ notably from those recorded in the Taipei Bear Vouchers 1.0 program. According to the Taipei City Government (2022), under the 1.0 program, the sports voucher generated the highest induced consumption effect, approximately 100%, followed by the dining voucher at around 79%. In contrast, the market, cultural, and accommodation vouchers exhibited relatively lower induced consumption rates, ranging from 11% to 40%. [^6] Taipei City Government (2022) does not report induced consumption effect for the Taipei Bear Vouchers 1.0 program disaggregated by demographic characteristics.
5.4 Intensity of Treatment Effect
@@
5.5 Comparison to Previous Studies
Similar to Xing et al. (2023), this paper examines consumer behavioral responses and the policy effectiveness of digital consumption vouchers. Xing et al. (2023) reported that China’s digital voucher policy generated an average of 3.07 yuan in consumer spending for every 1 yuan of government subsidy. This induced consumption effect is substantially stronger than that observed under Taipei Bear Vouchers 2.0 program. A plausible explanation is that the Chinese policy imposed a short, seven-day usage period, which may have heightened consumers’ sense of urgency and encouraged more immediate spending behavior.
Further insights can be drawn by comparing the results to Taiwan’s 2009 paper-based voucher program. According to Kan et al. (2017), the average MPC under the 2009 policy was approximately 0.25, indicating that each NT\$1 distributed in vouchers led to about NT\$0.24 in unplanned spending. The study also found that consumers who used the vouchers primarily for essential goods exhibited significantly lower MPC and contributed relatively little additional out-of-pocket spending. In contrast, individuals who were younger, more highly educated, more optimistic about the economic outlook, had higher household income, received larger voucher amounts, or experienced recent income growth tended to exhibit higher MPC. Overall, most of the 2009 paper-based vouchers were used for planned purchases. For example, nearly 90% of spending on essential goods was classified as substitutional, with only 10.6% attributable to the induced consumption effect. Spending on durable goods and services had an estimated substitution rate of around 75%.
In contrast, the dining voucher under the Taipei Bear Vouchers 2.0 program, also related to essential consumption, exhibited a significantly lower substitution rate, ranging between 23% and 38%. This suggests stronger performance in reducing substitution effects relative to the 2009 program. Moreover, the present study finds that the induced consumption rates for all types of Taipei Bear Vouchers 2.0 exceeded 70%, indicating a relatively high level of effectiveness in stimulating additional consumer spending.
Differences in consumer responses across the 2009 paper-based voucher program and the Taipei Bear Vouchers 2.0 program may reflect both variations in policy design and underlying consumer characteristics. The 2009 program was implemented nationwide, with participants geographically dispersed and permitted to spend vouchers across a broad range of categories without restriction. In contrast, the Taipei Bear Vouchers 2.0 program was limited to designated spending categories within Taipei City and targeted a more localized population. These structural differences likely shaped consumer spending behavior and, in turn, influenced the effectiveness of the two programs. Another relevant consideration is that our analysis focuses on the treatment effect among individuals who registered for the digital vouchers (i.e., the treated group), and therefore does not account for those who chose not to register (the latent group). If we were to estimate the unconditional treatment effects, i.e., the effects of distributing vouchers to all eligible individuals, the behavioral responses of the latent group could potentially alter the findings. However, given the specific policy context and the objective of assessing the realized impact of the program, our study deliberately concentrates on the treated group, which seeks to measure the effects experienced by voucher registrants and the consequent industry benefits.
@@ 5.6 Economic Impacts of Consumption Vouchers
To assess the economic impact of Taipei’s consumption voucher program, we apply the Taipei Input–Output model to evaluate the effects of the Taipei Bear Vouchers 2.0 policy and to calculate the associated output multipliers based on Equation (8). In this framework, the estimated total induced consumption is treated as a change in final demand, denoted by $\Delta\mathbf{F}$. Specifically, for each voucher type $k$, we compute the induced consumption as the product of the original issued amount and the behavioral adjustment factor $(1-ES_{k})\times(1+IC_{k})$, where $ES_{k}$ denotes the expenditure substitution rate and $IC_{k}$ represents the induced consumption rate. These adjusted values are then used as inputs in the input–output model to estimate the broader economic impacts. The coefficient matrix and value-added vector used in Equation (8) are provided in Table 7 in the Appendix.
@@
Table 5: Estimated Policy Inputs for the Taipei Bear Vouchers 2.0 Program
@@
Note: The original issued amounts are provided by the Taipei City Government, and these figures include the second-round extra stimulus vouchers. The optimistic and pessimistic estimates of the induced change in demand are calculated using the expression $(1-ES_{k})\times(1+IC_{k})$, where $ES_{k}$ denotes the substitution effect and $IC_{k}$ denotes the induced consumption effect for voucher type $k$. These estimates are obtained by multiplying the issued amount by the above expression, using the mean values from the lower and upper bounds of the respective confidence intervals for substitution and induced consumption effects. The monetary values in the table are expressed in 2022 price levels.
Table 6 presents the estimated economic benefits of the Taipei Bear Vouchers 2.0 program on Taipei City’s industrial economy under three scenarios. The baseline scenario reflects the economic impact based solely on the original budget allocation, treating it as a direct change in final demand without accounting for expenditure substitution or induced consumption effects. This scenario assumes a total policy input of NT\$584.53 million. In addition to the baseline, the table includes a pessimistic scenario and an optimistic scenario, both of which incorporate behaviorally adjusted estimates of final demand based on substitution and induced consumption effects as shown in Table 5.777Since the Taipei City input–output table used in this paper is based on the year 2016, the estimated final demand values for each type of Taipei Bear Vouchers 2.0 under different scenarios in 2022 were first deflated to 2016 price levels using the Consumer Price Index (CPI) before conducting the input–output analysis. Subsequently, the estimated economic benefits derived from the input–output analysis were re-inflated to 2022 monetary values using the CPI as the inflation adjustment index.
@@
Table 6: Economic Benefit of Taipei Bear Vouchers 2.0 Program
@@
Note: The output multiplier is defined as the ratio of the change in gross domestic product (GDP) to the original amount of policy expenditure. The original amount of policy expenditure is NT\$584.53 million. All units are in measured in million except the output multiplier. The monetary values in the table are expressed in 2022 price levels.